### v1

# CuLEmo: Cultural Lenses on Emotion - Benchmarking LLMs for Cross-Cultural Emotion Understanding

Abstract
NLP research has increasingly focused on subjective tasks such as emotion analysis. However, existing emotion benchmarks suffer from two major shortcomings: (1) they largely rely on keyword-based emotion recognition, overlooking crucial cultural dimensions required for deeper emotion understanding, and (2) many are created by translating English-annotated data into other languages, leading to potentially unreliable evaluation. To address these issues, we introduce Cultural Lenses on Emotion (CuLEmo), the first benchmark designed to evaluate culture-aware emotion prediction across six languages: Amharic, Arabic, English, German, Hindi, and Spanish. CuLEmo comprises 400 crafted questions per language, each requiring nuanced cultural reasoning and understanding. We use this benchmark to evaluate several state-of-the-art LLMs on culture-aware emotion prediction and sentiment analysis tasks. Our findings reveal that (1) emotion conceptualizations vary significantly across languages and cultures, (2) LLMs performance likewise varies by language and cultural context, and (3) prompting in English with explicit country context often outperforms in-language prompts for culture-aware emotion and sentiment understanding. We hope this benchmark guides future research toward developing more culturally aligned NLP systems.
CuLEmo: Cultural Lenses on Emotion - Benchmarking LLMs
for Cross-Cultural Emotion Understanding
Tadesse Destaw Belay1,2, Ahmed Haj Ahmed3, Alvin Grissom II3, Iqra Ameer4, Grigori Sidorov1, Olga Kolesnikova1, Seid Muhie Yimam5 1Instituto Politécnico Nacional, 3Haverford College, 3Wollo University, 4Pennsylvania State University, 5University of Hamburg, Contact: tadesseit@gmail.com
1 Introduction
Despite progress in bridging language barriers Ahuja et al. (2023), large language models (LLMs) still struggle to capture cultural nuances and adapt to specific cultural contexts Shen et al. (2024). Ideally, multilingual LLMs can not only facilitate cross-lingual communication but also incorporate an awareness of cultural sensitivities (i.e., what is deemed acceptable, normal, or inappropriate in a given culture), integrating such knowledge to foster deeper global connections Liu et al. (2024a).
Figure 1: CuLEmo creation pipeline and evaluations of LLMs in emotion and sentiment tasks.
LLMs and agent systems are being employed to interact extensively with humans across various applications such as customer service, healthcare, and education Wang et al. (2024). To facilitate effective interaction, incorporating aspects of cognitive and emotional-social intelligence—the ability to recognize and interpret human emotions Mathur et al. (2024)—can facilitate better interactions with more people. Factors such as age, cultural background, and personal experiences influence how individuals perceive and process information, particularly within subjective NLP tasks. Among these, emotion recognition (ER) and sentiment analysis (SA) are particularly sensitive to language- and culture-specific nuances Plaza-del-Arco et al. (2024).
Natural language frequently encodes emotional information Jim et al. (2024). For example, consider tipping customs in restaurants: in some cultures (e.g. North America), tipping is widely practiced, whereas in China, it is rare, and in Japan, it may even be considered offensive Givi and Galak (2017). Such differences underscore the importance of culturally-aware language technologies.
Although prior work has attempted cross-lingual emotion evaluation by translating emotion-annotated data in English into other languages Tahir et al. (2023); De Bruyne (2023), relying solely on translations from English can introduce incomplete or misleading insights. A more fair comparison requires the same underlying scenarios, each annotated natively across different languages and cultures. While emotion is language- and culture-dependent Plaza-del-Arco et al. (2024), comprehensive cross-cultural evaluations remain largely unexplored.
To bridge this gap, we propose Cultural Emotion (CuLEmo), a novel dataset that captures events and annotates them across multiple cultures and languages from scratch. CuLEmo enables the evaluation of multilingual LLM performance in analogous scenarios across different cultural contexts. Figure 1 shows the CuLEmo dataset creation and evaluation pipeline.
Culture can manifest in 1) the language of the data itself and 2) the annotation labels (i.e., multi-culturally informed annotations) Liu et al. (2024b). CuLEmo satisfies both conditions: it is multilingual and includes culturally grounded annotations. Indeed, the same event may evoke distinct emotional reactions in different cultures. In light of this, we pose the following research questions (RQs):
  1. 1.
    Do LLMs provide culturally aware-emotional responses?
  2. 2.
    Which cultures are more effectively represented in LLMs?
  3. 3.
    Can LLMs identify a country’s culture based on the text describing an event on the prompt used?
  4. 4.
    Does the language of the prompt affect an LLM’s ability for culture-aware emotion understanding?
To that end, this paper makes three key contributions. First, we introduce CuLEmo, a high-quality, multicultural, and multilingual benchmark dataset. Second, we leverage CuLEmo to investigate whether widely used multilingual LLMs can capture variations in emotional expression across cultures and languages in emotion and sentiment tasks. Finally, we highlight the variation in performance on culture-aware emotion understanding when LLMs are prompted in different languages.
2 Related Work
We now review related work in culturally-aware NLP, culture-oriented benchmarks, and cross-lingual study of linguistic emotional expression. Although culture is a complex concept, most definitions of culture encompass people, groups of people, and interactions between individuals and groups Liu et al. (2024c). In NLP, understanding culture is important for the safety and fairness of Large Language Models (LLMs).
2.1 Culture-oriented Benchmarks
Given the significance of culture in language model evaluation Adilazuarda et al. (2024), researchers have proposed various culture-oriented benchmarks to explore its effects on language understanding and generation. These efforts typically involve collecting and annotating multilingual and multicultural corpora to study cultural-driven phenomena in downstream NLP tasks. For instance, prior work has examined cross-cultural user statements Liu et al. (2021); Nayak et al. (2024), detected cultural differences and user attributes Sweed and Shahaf (2021); Qian et al. (2021), studied multilingual moral understanding Guan et al. (2022); Mohamed et al. (2022), and addressed culture-specific time expression grounding Shwartz (2022); Fung et al. (2023).
2.2 Emotion Across Languages
Several studies have examined multilingual emotion datasets. Some findings suggest that emotion categories can be preserved through machine translation, for example, by exploring how English-annotated emotion data translate into Finnish, French, German, Hindi, and Italian Kajava et al. (2020); Tahir et al. (2023); Bianchi et al. (2022). This work suggests that the changes in emotion labels often stem from the inherent difficulty of annotation rather than from linguistic differences.
Conversely, other works emphasize that emotions may not be consistently preserved across different languages. De Bruyne et al. (2022) showed that typologically dissimilar languages pose challenges for cross-lingual learning with mBERT-based models. De Bruyne (2023) argued that translation could fail to capture language-specific verbalizations and connotations—especially if certain emotion keywords do not exist in a given language (e.g., there is no direct word for “sadness” in Tahitian, and Amharic does not have an exact term for “surprise”). Qian et al. (2023) found that roughly half of the machine-translated outputs from English to Chinese fail to adequately preserve the original emotion, attributing these discrepancies to emotion-specific words and complex linguistic phenomena.
2.3 Emotions and Cultures
Recent work examines how cultural contexts shape emotional expression across languages. Havaldar et al. (2023a) analyzed embeddings of 271 emotion keywords in English, Spanish, Chinese, and Japanese by projecting into a Valence-Arousal plane using XLM-RoBERTa Reimers and Gurevych (2020) embeddings, revealing that multilingual models embed non-English emotion words differently. Havaldar et al. (2023b) analyzes Pride/Shame as a known cultural difference by prompting GPT-3.5 and GPT-4 to explore how these models handle pride and shame in the USA vs. Japan. They found that GPT-3.5 displayed limited knowledge of culturally specific norms. Ahmad et al. (2024) expands the 19 cultural questions of Havaldar et al. (2023b) to 37 questions and evaluates ChatGPT for the low-resource Hausa language for sentiment analysis. However, these evaluations are limited to 19/37 culturally relevant questions (not diverse and representative data), limited classes (Pride/Shame or positive/negative/mixed), and limited cultures (USA vs. Japan or Hausa).
3 CuLEmo Dataset Preparation
We now describe the precise steps in creating the CuLEmo dataset.
3.1 Collecting Cultural Events
We manually crafted scenarios, searched online, and prompted LLMs to gather traditions, events, norms, and actions that elicit culturally different emotions across six target countries (USA, UAE, Germany, Ethiopia, India, and Mexico).We drew inspiration from the work of Havaldar et al. (2023b) to enhance topic diversity. Language representatives are asked to propose events distinct from those of other countries in the form of emotion-oriented questions. Importantly, these events do not contain explicit emotional keywords (as typically seen in traditional emotion datasets). We also referred to the International Survey on Emotion Antecedents and Reactions (ISEAR) (Scherer and Wallbott, 1994) data format, "When I … situations that cause a specific emotion"—a well-known English dataset for emotion analysis consisting of self-reported events from around 3,000 respondents across 37 countries and five continents. Table 1 displays the ten broad categories and the number of questions in each category.
Categories # Qn.
Family relationships 45
Social etiquette and interactions 65
Personal appearance and dress code 32
Cultural and religious practices 62
Sexual and intimate relationships 38
Professional contexts 28
Food and dining etiquette 35
Personal privacy 25
Emotional and psychological situations 40
Public behavior and norms 30
Total 400
Table 1: Ten broad categories and the number of events (Qn - questions) in the CuLEmo dataset.
3.2 Human-Adapted Translation
After collecting the events in English, we translated them into five other target languages—Arabic, Amharic, German, Hindi, and Spanish—using Google Translate, followed by native-speaker approvals. Because the questions are simple “How do you feel when … ?” and lack explicit emotion keywords, translation quality did not affect their cultural content. While we acknowledge that other existing works often depend on translating emotion-annotated data from English into other languages with its labels Kajava et al. (2020); Tahir et al. (2023); Bianchi et al. (2022), the translation process we employed in our study is done before any annotation. To ensure correctness, native speakers evaluated the translations. While most translations were acceptable, a few adjustments were made—for example, for Amharic to fix gender references and timing expressions.
3.3 Language and Cultures Covered
Several factors guided our choice of languages: (1) typological variety (five distinct languages with four different scripts), (2) geographical diversity (eastern vs. western contexts), (3) resource availability (low- vs. high-resource languages), and (4) the availability of native speakers for translation reviews. These considerations helped ensure a broad range of cultural norms and concepts.
3.4 CuLEmo Annotation
We used Amazon Mechanical Turk (MTurk) for most annotations, ensuring at least five native-speaker annotators per instance from the respective targeted country. We used a customized POTATO Pei et al. (2022) annotation tool for languages lacking sufficient MTurk annotators (e.g., Amharic) and recruited local native speakers who met our criteria. Annotators were fairly compensated \$12/hr. (Prolific [^1] https://www.prolific.com/’s minimum annotation wage is \$9/hr). In cases where no majority vote emerged among five annotators, we collected two additional annotations and used majority voting to finalize the label.
Figure 2: Emotion label distribution across countries/languages.
Figure 2 illustrates the distribution of emotion labels. We use six categories: joy, fear, sadness, anger, guilt, and neutral (no specific emotion). These labels were adapted from De Bruyne et al. (2019), where five categories were clustered, plus a neutral class. Our annotation guidelines group related labels as “helper” categories—for instance, “love” and “happy” under joy, and “shame” under guilt—to assist annotators in selecting the most appropriate coarse-grained label. Examples from the dataset are provided in Appendix A.
Pairwise Label Agreements Across Countries. We further examined label differences across countries by calculating pairwise agreements after majority voting (Figure 3). Ethiopia (ET) and the United Arab Emirates (AE) exhibit the highest agreement at 55%, along with Germany (DE) and the United Arab Emirates (AE), both exhibit a high neutral class; see Figure 2. In contrast, Germany (DE) and India (IN) show the lowest agreement, at 29.0%. Labels from India thus tend to diverge from those of other countries.
Figure 3: Pairwise emotion label agreements across countries (ordered by their average agreement). Abbreviations: US = USA, MX = Mexico, DE = Germany, AE = UAE, ET = Ethiopia, and IN = India.
4 Experimental Setup
4.1 Task Formulation
To investigate the emotional comprehension capabilities of LLMs, we examine how they associate different countries’ cultures with their corresponding languages. Specifically, we explore culture-aware emotion understanding via two main tasks: emotion prediction and sentiment analysis. All tested models are instruction-fine-tuned, except for the Aya-expanse model. We also experiment with prompts that do and do not include explicit country context, using the phrase "You live in <<country name>>," (where <<country name>> is one of the six targeted countries: USA, UAE, Ethiopia, Germany, India, and Mexico). Each task is framed as a text generation problem, and the models are evaluated in a zero-shot setting.
4.2 Model Selection
We evaluate a variety of recent LLMs known for strong performance on standard benchmarks. We aim to include both smaller and medium-sized models, as well as open-source and proprietary options. The models evaluated include:
  1. 1.
    Open-Source Models: LLAMA-3 (3.2-3B, 3.1-8B) Dubey et al. (2024); Meta AI (2024), GEMMA (2B, 9B) Team Gemma et al. (2024), Aya (expanse-8b, 101-13B) Üstün et al. (2024); Dang et al. (2024), Ministral (3B, 8B) Mistral AI (2024)
  2. 2.
    Proprietary Models: GPT (3.5, 4) OpenAI (2024), GEMINI-1.5 Google Deepmind (2024), Claude (3.5-sonnet, 3-opus) Anthropic AI (2024)
4.3 Multilingual Prompt Construction
To examine the impact of prompt language on model performance—and to assess each model’s cultural awareness—we design both English and in-language prompts. The instruction, input text, and expected answer are in English when using English prompts. For in-language prompts, all elements are in one of the five target languages—Arabic (AR), German (DE), Amharic (AM), Hindi (HI), or Spanish (ES). Complete examples of our multilingual prompts for both emotion prediction and sentiment analysis are provided in Appendix C. We extract each model’s answer from its generated text using PEDANTS Li et al. (2024).
5 Result and Analysis
5.1 Culture-Aware Emotion Prediction
Do LLMs provide culturally aware emotional responses? A culture-aware model should accurately answer questions related to any culture, demonstrating uniformly high accuracy. To test this, we assess each LLM’s emotional understanding using English and in-language prompts with the context "You live in <<country name>>,".
LLMs USA UAE Germany Ethiopia India Mexico
EN EN AR EN DE EN AM EN HI EN ES
Llama-3.2-3B 0.44 0.20 0.12 0.20 0.20 0.18 0.30 0.26 0.26 0.28 0.37
Llama-3.1-8B 0.58 0.52 0.47 0.48 0.39 0.48 0.14 0.37 0.34 0.56 0.57
Gemma-2-2B 0.62 0.59 0.45 0.64 0.54 0.48 0.17 0.38 0.29 0.58 0.56
Gemma-2-9B 0.57 0.51 0.56 0.47 0.53 0.47 0.29 0.40 0.34 0.53 0.60
Aya-expanse-8b* 0.38 0.28 0.20 0.30 0.24 0.30 0.29 0.32 0.29 0.37 0.47
Aya-101-13B 0.60 0.60 0.43 0.68 0.43 0.52 0.34 0.39 0.34 0.56 0.48
Ministral-8B 0.65 0.61 0.06 0.58 0.72 0.49 0.23 0.39 0.19 0.57 0.32
Claude-3.5-sonnet 0.57 0.48 0.54 0.46 0.42 0.51 0.49 0.40 0.36 0.58 0.61
Claude-3-opus 0.54 0.48 0.47 0.43 0.32 0.53 0.43 0.37 0.36 0.61 0.61
Gemini1.5-flash 0.56 0.56 0.56 0.46 0.48 0.51 0.51 0.41 0.41 0.62 0.64
GPT-4 0.60 0.55 0.48 0.51 0.50 0.54 0.29 0.40 0.40 0.64 0.65
Table 2: LLMs’ accuracy for the emotion prediction task. Columns labeled EN/AR/DE/AM/HI/ES show the prompt language for each corresponding culture (e.g., UAE is tested with English and Arabic). The highest-scoring model across English and in-language prompts is highlighted in bold, and the best model for the in-language prompt is underline. * is not an instruction fine-tuned model.
Results: Table 2 presents the accuracy of each LLM for culture-aware emotion prediction. The choice of prompt language significantly influences performance. Generally, proprietary models are less affected by in-language prompts compared to open-source models, especially for Spanish and German. Certain cultures appear better represented in the models—Ministral-8B scores highly on German (72%), and GPT-4 performs best on Mexican (65%). In contrast, performance in Indian culture (Hindi) lags, particularly those using Hindi, Amharic, or Arabic scripts. Larger models do not always outperform smaller ones; Gemma-2-2B and Ministral-8B show competitive or superior accuracy relative to some proprietary models. When prompted in English, all models achieve a reasonable accuracy. Ministral-8B can exceed proprietary performance in English and German. GPT-4 answers the same emotion while using the corresponding country name in the prompting; see predicted examples in Table 6. Overall, results suggest that culture-aware emotion understanding remains challenging for the tested LLMs, especially for lesser-resourced languages and cultures.
5.2 Culture Representation in LLMs
Which culture is more represented in LLMs? Here, we evaluate LLM performance using English prompts without any explicit country context. We then measure how models respond to events from each target country.
LLMs Country
US AE DE ET IN MX
Llama-3.2-3B 0.44 0.41 0.36 0.35 0.31 0.46
Llama-3.1-8B 0.62 0.54 0.50 0.51 0.36 0.60
Gemma-2-2B 0.61 0.59 0.68 0.49 0.32 0.57
Gemma-2-9B 0.59 0.54 0.54 0.50 0.41 0.57
Aya-101-13B 0.61 0.57 0.60 0.55 0.38 0.58
Ministral-8B 0.66 0.60 0.56 0.50 0.38 0.59
Claude-3.5-s. 0.56 0.51 0.39 0.49 0.36 0.60
Claude-3-op 0.52 0.40 0.37 0.49 0.36 0.59
Gemini1.5 0.53 0.51 0.39 0.48 0.39 0.59
GPT-4 0.58 0.55 0.46 0.48 0.37 0.63
Table 3: LLMs’ accuracy using English prompts without specifying the country context, "You live in «country name»". The columns list each target country. The bold columns indicate the top three most accurately represented cultures.
Results: According to Table 3, the USA, Mexico, and Germany consistently achieve higher accuracy scores, while the UAE, Ethiopia, and India remain less accurately represented. This suggests that certain cultures may be more prevalent in the underlying training data.
5.3 Does Language Represent Country?
Can LLMs identify one country’s culture based solely on the prompt language? In this experiment, we remove explicit country context (e.g., "You live in <<country name>>") and test if LLMs can infer cultural cues only from the language used in the prompt.
LLMs Language
EN AR DE AM HI ES
Llama-3.2-3B 0.44 0.13 0.19 0.30 0.25 0.37
Llama-3.1-8B 0.62 0.46 0.31 0.18 0.32 0.56
Gemma-2-2B 0.61 0.47 0.56 0.14 0.27 0.52
Gemma-2-9B 0.59 0.59 0.59 0.28 0.34 0.61
Aya-101-13B 0.61 0.34 0.39 0.31 0.30 0.49
Ministral-8B 0.66 0.07 0.67 0.20 0.23 0.34
Claude-3.5-s. 0.56 0.51 0.35 0.53 0.38 0.61
Claude-3-op 0.52 0.45 0.30 0.47 0.37 0.58
Gemini1.5 0.53 0.54 0.42 0.52 0.37 0.65
GPT-4 0.58 0.46 0.45 0.27 0.41 0.64
Table 4: LLMs’ accuracy for in-language prompts without country context. For each language, bold marks the best performance.
Results: Table 4 shows that accuracy drop significantly when country context is omitted. Comparing these scores with Table 2 (where country context is included), we see consistent performance boosts (e.g. +1% in Spanish with GPT-4, +5% in Arabic with Gemini1.5, +6% in German with Ministral, +9% in Amharic with Claude-3.5-sonnet, and +21% in Hindi with GPT-4) when the prompt explicitly names the country. This indicates that language alone does not reliably convey cultural context. Notably, models like Claude-3.5-sonnet and GPT-4 show improvements in Indian and Ethiopian data when country context is specified, while English prompts (without “USA” context) are less affected. Overall, providing country name information remains crucial for accurate culture-aware emotion understanding, especially for less-resourced languages.
5.4 Culture-Aware Sentiment Analysis
For the sentiment analysis experiment, we follow De Bruyne et al. (2019); Mostafazadeh Davani et al. (2022) by grouping emotions into positive (joy), negative (fear, anger, guilt, sadness), and neutral sentiments. Table 5 shows the accuracy of each LLM under this three-class setup.
LLMs USA UAE Germany Ethiopia India Mexico
EN EN AR EN DE EN AM EN HI EN ES
Llama-3.2-3B 0.57 0.41 0.15 0.20 0.24 0.48 0.51 0.60 0.56 0.55 0.60
Llama-3.1-8B 0.65 0.58 0.50 0.49 0.43 0.59 0.43 0.57 0.55 0.64 0.68
Gemma-2-2B 0.68 0.65 0.58 0.66 0.59 0.56 0.46 0.50 0.59 0.66 0.66
Gemma-2-9B 0.55 0.57 0.63 0.49 0.56 0.62 0.50 0.57 0.61 0.63 0.68
Aya-expanse-8b* 0.45 0.33 0.26 0.31 0.27 0.39 0.50 0.45 0.58 0.45 0.57
Aya-101-13B 0.65 0.64 0.55 0.69 0.45 0.55 0.53 0.52 0.56 0.61 0.63
Ministral-8B 0.70 0.66 0.06 0.60 0.73 0.59 0.46 0.54 0.24 0.65 0.53
Claude-3.5-sonnet 0.65 0.58 0.61 0.49 0.46 0.66 0.62 0.60 0.61 0.66 0.72
Claude-3-opus 0.63 0.60 0.56 0.46 0.35 0.67 0.58 0.59 0.61 0.72 0.75
Gemini1.5-flash 0.63 0.63 0.63 0.48 0.50 0.63 0.60 0.60 0.53 0.69 0.71
GPT-4 0.67 0.61 0.54 0.48 0.51 0.62 0.49 0.58 0.62 0.72 0.75
Table 5: LLMs’ accuracy for culture-aware sentiment analysis (positive/negative/neutral). Columns correspond to English (EN) prompts and the native language (e.g., Arabic for UAE). The highest-scoring model across English and in-language prompts is highlighted in bold, and the best model for the in-language prompt is underlined. * is not an instruction fine-tuned model.
Results: Table 5 shows better overall performance on sentiment analysis compared to fine-grained emotion classification. For example, GPT-4 gains notable accuracy (e.g., +22% for Hindi). Similarly, the highest score appears for Mexican culture (Spanish) with Claude-3-opus and GPT-4, each at 75%. By contrast, the non-instruction-tuned Aya-expanse-8b struggles more as it is not instruction fine-tuned. Smaller models like Gemma-2-2B remain competitive with proprietary systems. Nonetheless, performance drops persist for Amharic and Hindi, reflecting the continued challenges of culture-aware tasks in lower-resource contexts.
Questions US AE DE ET IN MX
How would you feel when you did not tip the waiter at the restaurant? guilt neutral neutral guilt guilt guilt
guilt guilt guilt guilt guilt guilt
How would you feel when someone insults someone’s religion? neutral anger neutral anger neutral anger
anger anger anger anger anger anger
How would you feel if someone wears black to a wedding? sadness neutral neutral anger neutral neutral
neutral neutral neutral neutral neutral neutral
How would you feel when you see a female wearing small pants on the street? neutral sadness neutral anger neutral neutral
neutral neutral neutral neutral neutral neutral
How would you feel when your attendee joined the meeting after 10 minutes started? anger neutral anger neutral anger anger
neutral anger anger anger anger anger
How would you feel if your parents arranged a marriage for you without your input? anger anger neutral neutral neutral neutral
anger anger anger anger anger anger
How would you feel upon receiving the message that you have been accepted as a medical student? joy joy neutral joy joy joy
joy joy joy joy joy joy
Table 6: Examples of predictions from GPT-4 model. The first row under the country-code columns in each section represents the gold-label emotions, while the second row displays the predicted label. The predictions are using country context, by adding explicitly You live <<country name>> in the prompt
6 Discussion
Our analyses provide several insights into the current state of LLMs with respect to cultural emotion understanding. We highlight three main lessons learned and propose potential steps to enhance the cultural awareness of LLMs.
6.1 Emotion Varies Across Languages and Cultures
Data analysis from the CuLEmo dataset in Figure 2 shows notable differences in how annotators from different countries perceive the same event. For instance, Germany has the highest proportion of neutral (no emotion) labels. Figure 4 further illustrates the distribution of positive, negative, and neutral sentiments across 400 instances in each country.
How are emotions distributed across languages and cultures? Based on the dataset analysis of emotion distribution across languages, shown in Figure 2: German (87%), Arabic (58%), and English (50.5%) data have the most neutral (no emotion). Amharic (29.5%), Arabic (22%), and Spanish (21.5%) languages have the most anger emotion. These findings confirm that a single event can evoke distinct emotional reactions depending on cultural background and language.
Figure 4: Sentiment distribution across countries in the CuLEmo dataset (400 total instances).
6.2 Prompt Language Strongly Affects Cultural Emotion Understanding
As illustrated in Table 2, prompt language plays a major role in LLM performance for emotion prediction and sentiment analysis. For less-resourced languages like Hindi and Amharic, prompting in English consistently yields better results—sometimes by as much as a 20% improvement. Conversely, in-language prompts with explicit country context tend to work best for high-resource languages such as German and Spanish. These discrepancies stem from differences in both linguistic coverage and instruction-following abilities learned during pre-training. One practical solution is to leverage English prompts while specifying the target country (e.g., “You live in <<country name>>,”).
6.3 Performance Gaps Reflect Under-Represented cultures
We observe notably lower accuracy for Ethiopia and India in both emotion and sentiment tasks, suggesting that models may be less familiar with cultural practices and norms for these under-represented contexts. Ensuring greater diversity in training corpora is key to improving model performance for such cultures. Providing explicit country context can partially offset these gaps by nudging models to incorporate relevant cultural knowledge.
Overall, our findings underscore the importance of cultural context in developing and deploying LLMs. Beyond balanced data collection, researchers may explore culture-specific tuning or reinforcement learning from human feedback to further refine models’ ability to interpret and respect cultural nuances.
7 Conclusion
In this paper, we evaluated a diverse set of state-of-the-art LLMs for their ability to predict culturally aware emotion prediction and sentiment analysis tasks. We investigated the influence of including explicit country references “You live in <<country name>>” and varying the query language. Our results indicate that LLMs tend to excel at culturally driven emotions that are well-represented in their training data and underperform for less represented cultures. Specifically, we found that 1) Emotion is culture-dependent and can vary notably across languages and regions, 2) LLMs exhibit sizable performance gaps when tested on culture-specific emotions from under-represented locales, 3) Providing explicit country context in prompts improves both emotion and sentiment prediction, and 4) Sentiment analysis is better for the models, likely because it involved few class (positive, negative, neutral) than fine-grained emotion categories.
Moving forward, we suggest training LLMs in approaches such as 1) enriching the training data with diverse cultural information from various sources like literature, news, and cultural databases and 2) implementing fine-tuning techniques that specifically train the LLM on prompts and datasets focused on different cultural contexts so that they can be culturally aware and able to generate responses that are sensitive to cultural nuances. We also encourage more extensive evaluation of multilingual models using benchmarks designed to measure cultural awareness alongside standard accuracy metrics. Future research could also explore the influence of annotator demographics—such as age, gender, education level, political stance, and religion—on culture-specific emotion annotation. Finally, we hope that releasing the CuLEmo dataset will foster further exploration into culturally nuanced NLP tasks and lead to more inclusive language models.
Limitations
Emotional subjectivity remains a central challenge in emotion analysis tasks. Although annotating data via crowdsourcing such as Amazon Mechanical Turk (MTurk) is common in NLP dataset creation Mohammad et al. (2018), and despite applying strict qualification criteria for annotators, maintaining consistent annotations is difficult given the inherently subjective nature of emotions.
Additional constraints arise from our decision to limit the emotion label space to six classes; including more categories (e.g. surprise or disgust) could yield more fine-grained insights Niu et al. (2024). Our dataset also covers only six languages/countries and comprises 400 events, which may restrict generalizability. Data bias is another concern, as emotions and cultures vary widely and cannot be exhaustively captured.
Regarding open-source LLMs, we opted to evaluate only small- (2B,3B) and medium-sized (8B, 9B, 13B) models due to resource constraints and for experimental reproducibility. While larger LLMs might achieve higher accuracies across target languages, they remain beyond the scope of our current setup. Finally, although 400 events allow for controlled experiments, evaluating models on more extensive and varied data would provide a clearer picture of their culture-aware performance.
Appendix A Dataset Examples
The below box shows English examples from CuLEmo datasets.
English Examples from the CuLEmo dataset 1. How would you feel when you did not tip the waiter at the restaurant? 2. How would you feel when someone insults someone’s religion? 3. How would you feel if someone wear black to a wedding? 4. How would you feel when you see a female wearing small pants on the street 5. How would you feel when your attendee joined the meeting after 10 minutes started 6. How would you feel when you saw a young dog running to you? 7. How would you feel when someone accepted your gift in one hand? 8. How would you feel when you crossed an older man in the street? 9. How would you feel when someone calls a religious leader by their name 10. You are a doctor. How would you feel when someone calls you by your name without the doctor prefix 11. How would you feel if your parents arranged a marriage for you without your input? 12. How would you feel if someone touched or shook hands with you while you were of the opposite sex? 13. How would you feel when a guest leaves food on their plate at your home? 14. How would you feel if you saw a woman not cover her hair while entering a religious place? 15. How would you feel when you were expected to wash other legs 16. How would you feel if someone refused your gift before accepting it? 17. How would you feel when you kissed your girlfriend/boyfriend and saw your father? 18. How would you feel when your parent caught you with a pornographic film? 19. If someone is passing near to you, how would you feel when they didn’t give you a greeting? 20. How would you feel when you see a female eating on the street? 21. How would you feel when you see a Sheikh drinking alcohol? 22. How would you feel when you see a priest drinking alcohol? 23. How would you feel when you see homosexual friends? 24. How would you feel when someone shows you his middle finger gesture? 25. How would you feel when someone gave the money in their left hand? 26. How would you feel if your guests chose to keep their shoes on when entering your home? 27. How would you feel if a friend offered to pay for your meal in a restaurant? 28. How would you feel if your child expressed a desire to marry someone from a different religion? 29. How would you feel if someone bargained for your prices? 30. When you are at a meal, how would you feel about being told to wait for the host to say start before starting to eat? 31. How would you feel when you expected to wait for the eldest person to start eating before you begin? 32. How would you feel if someone were scrolling their phone during a meal? 33. How would you feel when you saw a young dog running toward you? 34. How would you feel when someone cheek-to-cheek kisses your wife/husband? 35. How would you feel if someone bargained for your prices? 36. How do you feel when a guest arrives late to your lunch invitation? 37. How do you feel seeing someone eating in hand without using utensils? 38. How would you feel if someone called an elder by their first name without a title? 39. How would you feel if a guest left your home without eating anything? 40. How would you feel if your friend didn’t stand up when an elder entered the room?
Appendix B Implementation Details
B.1 Hyper-parameters
For open-source LLMs, we used the default generation hyper-parameters (top-p sampling with p = 0.9 and temperature = 0.0, max_new_tokens=200). For others, we directly employed their pre-defined interfaces, either through their online API or the CHAT function from the Transformers library. The proprietary models are called through their API in Python. All open-source models are evaluated using default parameters from the hugging face.
B.2 LLM Versions
Versions of the proprietary LLMs and hugging face names for the open-source LLMs are given below.
  • Llama-3.2-3B Meta AI (2024) — meta-llama/Llama-3.2-3B-Instruct
  • Llama-3.1-8B Dubey et al. (2024) — meta-llama/Llama-3.1-8B-Instruct
  • Gemma-2-2B Team Gemma et al. (2024) — google/gemma-2-2b-it
  • Gemma-2-9B Team Gemma et al. (2024) — google/gemma-2-9b
  • Aya-expanse-8b Dang et al. (2024) — CohereForAI/aya-expanse-8b
  • Aya-101 Üstün et al. (2024) — CohereForAI/aya-101
  • Ministral-8B Mistral AI (2024) — mistralai/Ministral-8B-Instruct-2410
  • Claude-3.5-sonnet Anthropic AI (2024) — claude-3-5-sonnet-20240620
  • Claude-3-opus Anthropic AI (2024) — claude-3-opus-20240229
  • Gemini1.5-flash Google Deepmind (2024) — gemini-1.5-flash-002
  • GPT-4 OpenAI (2024) — gpt-4 (turbo-2024-04-09)
Appendix C Prompts
Table 7: Prompts used for probing emotions from LLMs. In English prompt, <<Country name>> will change accordingly from lists of countries [United States of America (USA), Ethiopia, United Arab Emirates (UAE), Germany, India, Mexico] based on the culture prompting. We enforce the model in the prompt to answer only one of the given options.
Appendix D English Prompt Results
Figure 5: English prompting results with You live in <<country name>> context for the corresponding languages and LLMs.
Appendix E MTurk Annotation Qualification Settings
To target suitable workers on MTurk, we set the following qualifications:
  1. 1.
    Location must be in the target country for each language by assuming annotators that live in the specified country are native or adopted the culture.
  2. 2.
    Number of HITs approved must exceed 1,000 to ensure experienced workers.
  3. 3.
    HIT approval rate must be at least 99%, favoring high-quality, consistent annotators